### v1

# Cuff-less Arterial Blood Pressure Waveform Synthesis from Single-site PPG using Transformer & Frequency-domain Learning

Abstract
We propose two novel purpose-built deep learning (DL) models for synthesis of the arterial blood pressure (ABP) waveform in a cuff-less manner, using a single-site photoplethysmography (PPG) signal. We utilize the public UCI dataset on cuff-less blood pressure (CLBP) estimation to train and evaluate our DL models. Firstly, we implement a transformer model that incorporates positional encoding, multi-head attention, layer normalization, and dropout techniques, and synthesizes the ABP waveform with a mean absolute error (MAE) of 14. Secondly, we implement a frequency-domain (FD) learning approach where we first obtain the discrete cosine transform (DCT) coefficients of the PPG and ABP signals corresponding to two cardiac cycles, and then learn a linear/non-linear (L/NL) regression between them. We learn that the FD L/NL regression model outperforms the transformer model by achieving an MAE of 11.87 and 8.01, for diastolic blood pressure (DBP) and systolic blood pressure (SBP), respectively. Our FD L/NL regression model also fulfills the AAMI criterion of utilizing data from more than 85 subjects, and achieves grade B by the BHS criterion.
Index Terms:
PPG, arterial blood pressure, systolic, diastolic, transformer, ridge regression, discrete cosine transform.
I Introduction
Hypertension (high blood pressure) and hypotension (low blood pressure) are two common blood pressure (BP) irregularities that are of great clinical significance. According to the World Health Organization (WHO), approximately 1.28 billion people aged 30-79 years suffer from hypertension, with 46% unaware of it and only 42% receiving treatment [1]. Further, hypertension is more prevalent in low and middle income countries [2]. Hypertension (also known as the silent killer) is a major risk factor for a range of cardiovascular diseases, stroke, chronic kidney disease and pre-mature death [3]. Hypertension, if not treated, could elevate both the systolic BP (SBP) and diastolic BP (DBP) from normal range (below 120/below 80) to stage 1 hypertension (130-139/80-89), to stage 2 hypertension (above 140/above 90), to hypertensive crisis (above 180/above 120), leading to health emergency necessitating immediate hospitalization. Hypotension, though less frequent than hypertension, is also widespread as approximately (10-20)of people of age over 65 years experience hypotension. Hypotension occurs when blood pressure falls below 90/60, which could lead to dizziness and nausea. Hypotension becomes perilous when BP falls below 60/45, causing blurriness, confusion, coma, or even death [4]. Thus, regular and frequent monitoring of blood pressure by the elderly is of utmost importance for early diagnosis and efficient management of hypertension and hypotension [5].
The existing clinical-grade methods for blood pressure estimation (including ambulatory devices) could be classified into two categories: i) invasive methods, ii) non-invasive but cuff-based methods. Invasive methods are the gold standard methods which measure the centralized aortic blood pressure by inserting catheters in the aorta (main artery near the heart) [6]. Need not to say that this method is painful, prone to infections, and required trained professionals in a clinical setting. Next, non-invasive but cuff-based methods work by inflating the cuff around the arm and rely upon manual auscultation and Sphygmomanometers for peripheral blood pressure estimation. These methods, though highly accurate, suffer from a number of limitations, e.g., inconvenience of regular monitoring, requirement of trained personnel, Sphygmomanometers use toxic mercury, etc [7, 8, 9]. Further, the need for frequent cuff inflation and deflation makes continuous BP measurement impractical, especially for individuals with weaker arms, or the elderly [10]. Oscillometric devices, though do not require manual auscultation, have relatively lower accuracy as they rely on mean arterial blood pressure (MABP) for estimating both SBP and DBP.
The limitations of the existing invasive methods and cuff-based non-invasive methods have thus prompted the researchers to design novel cuff-less blood pressure (CLBP) estimation methods which are meant to be user-friendly, reliable, and support continuous and at-home BP measurement [7, 11, 12, 13]. Additionally, advancements in wearable technology (e.g., smart watches, smart bands, fitness trackers, internet of medical things, flexible electronic tattoos, etc.) and assistive gadgets (e.g., smartphones, airpods, etc.) have further accelerated the development of the CLBP estimation methods [14], [15]. CLBP estimation methods either rely upon bio-signals such as photoplethysmography (PPG), electrocardiography (ECG), phonocardiography (PCG) acquired non-invasively from one or more body sites [12], or innovative sensor technology, e.g., tonometry [16].
Contributions. Inline with the recent works on CLBP estimation, we propose two novel purpose-built deep learning (DL) models for synthesis of the arterial blood pressure (ABP) waveform in a cuff-less manner, using a single-site PPG signal. We utilize the public UCI dataset on CLBP estimation to train and evaluate our DL models. The main contribution of this work is two-fold:
  • We implement a transformer model that incorporates positional encoding, multi-head attention, layer normalization, and dropout techniques, and synthesizes the ABP waveform with a mean absolute error (MAE) of 14.
  • We implement a frequency-domain (FD) learning approach where we first obtain the discrete cosine transform (DCT) coefficients of the PPG and ABP signals corresponding to two cardiac cycles, and then learn a linear/non-linear (L/NL) regression between them.
The transformer model (the FD L/NL regression model) achieves an MAE of 19 and 9 (11.87 and 8.01), for diastolic blood pressure (DBP) and systolic blood pressure (SBP), respectively. Our FD L/NL model achieves grade B by the British Hypertension Society (BHS) criterion.
Feasibility. PPG and ABP signals, though of different origins, have a very similar morphology (because both capture the pressure exerted by the blood volume on arterial walls). In other words, the morphology of the two signals is tightly binded to each other, at the sub-cardiac cycle resolution. Furthermore, the recent advances in the generative artificial intelligence (AI) has made it possible to translate one biosignal to another, e.g., [17] translates a PPG signal to an ECG signal, [18] translates a PPG signal to an ABP signal. This motivates us to propose purpose-built DL models for synthesis of the ABP waveform from a single-site PPG waveform.
Outline. Section II provides some necessary background and summarizes selected related work. Section III provides compact details of the public UCI dataset that we have used, and the key data pre-processing steps that we have performed. Section IV describes the two deep learning models we have implemented for cuff-less arterial blood pressure waveform synthesis. Section V presents some selected results. Section VI concludes the paper.
II Background & Related Work
II-A Background
Pulse wave analysis methods: During the last decade, researchers have made extensive investigation of the physiological waveforms acquired in a non-invasive manner (e.g., ECG, PPG, PCG, etc.), under the umbrella term of pulse wave analysis (PWA) [12]. The aim of PWA methods is to develop a number of indices to indirectly measure the blood pressure in a cuff-less manner using wearables (e.g., smart watches) and assistive gadgets (e.g., smartphones) [14]. Some prominent indices (markers for blood pressure) include the following: pulse transit time (PTT), pulse arrival time (PAT), and pulse wave velocity (PWV). Before we delve into the discussion of the related work, it is imperative to formally define the three indices (BP surrogates) as follows:
  • PTT is the time taken by a pulse pressure wave to travel from one point of an artery to another. PTT is inversely proportional to the blood pressure. A typical measurement system for PTT includes two PPG sensors placed at two different locations close to the artery [19], [20].
  • PAT is the time delay between the R-peak in the ECG signal and the time of arrival of the corresponding (PPG) pulse at the peripheral site, during the same heart beat. In other words, PAT is the time taken by the pulse pressure wave to travel from the heart to the peripheral measurement site. PAT is related to PTT as follows: PAT = PTT + pre-ejection period [12].
  • PWV is the velocity of a pulse pressure wave that travels a distance $L$ from a proximal point to a distal point. PWV is related to PTT as follows: $PWV=\frac{L}{PTT}$ [12].
Despite the good correlation of PTT, PAT and PWV with the blood pressure, these indices have a number of limitations as well, e.g., need for simultaneous and synchronous measurement of two waveforms from two different sites, calibration issues, lack of consensus on translation of PTT/PAT/PWV to BP (i.e., there exist a number of equations to map PTT/PAT/PWV to BP, each with many unknowns that are subject-specific), etc. [21], [12], [19]. This highlights the need for the design of novel single-site methods for cuff-less blood pressure (CLBP) estimation.
Regulatory criteria for CLBP estimation methods: Over time, two regulatory bodies have developed standards to grade the performance of the emerging CLBP estimation methods: 1) association for the advancement of medical instrumentation (AAMI), and 2) British hypertension society (BHS) [^1] Additionally, IEEE has also developed 1708-2014 (and 1708a-2019) standard for wearable cuffless blood pressure measuring devices [22]..
  • The AAMI requires a dataset consisting of at least 85 individuals having low, high, and normal blood pressure. Further, according to AAMI, the mean error (and standard deviation of the error) of a CLBP estimation method should be within 5 mmHg (and 8 mmHg) [23].
  • The BHS assigns one of the three letter grades (A, B, C) to each CLBP estimation method. If the model error is less than or equal to 5 mmHg 60% of time, 10 mmHg 85% of time, and 15 mmHg 95% of time, then the model is given grade A. Similarly, if the model error is less than 5 mmHg 50% of time, 10 mmHg 75% of time, and 15 mmHg 90% of time, then the model is given grade B. Finally, if the model error is less than 5 mmHg 40% of time, 10 mmHg 65% of time, and 15 mmHg 85% of time, then the model is given grade C [24].
II-B Related work
Since the literature on CLBP estimation is quite vast, we could only summarize selected related works due to space constraints (see [12], [19] and the references therein, which provide a more detailed treatment of the subject). The CLBP estimation methods are mostly PWA methods which are either two-sites/two-waveforms methods or single-site methods. More recently, researchers have extensively used various deep learning (DL) models for ABP waveform synthesis and CLBP estimation.
Single-site-based methods: [25] utilizes single-site PPG data collected from 150 subjects, computes PWV, and feeds it to a learning-based non-parametric regression method in order to estimate MABP, SBP and DBP. [26] collects multi-wavelength, single-site PPG data from 10 subjects, computes the PTT using the IR-PPG and blue-PPG, in order to estimate the DBP and SBP. [27] recruits 22 young healthy subjects to measure their ballistocardiogram (BCG) using the foot-based pressure sensor/force plate, in order to estimate the DBP (SBP) using I-J interval (J-K amplitude) of a BCG. [18] utilizes the UCI dataset and implements the well-known U-Net model to synthesize an ABP waveform from a single PPG waveform.
Two-sites-based methods: [28] proposes a two-sites-based method for CLBP estimation that utilizes a PPG sensor at the fingertip and an impedance plethysmography (IPG) sensor at the wrist in order to compute the PTT which is then translated to SBP and DBP values. [29] recruits 12 young healthy subjects, records their ECG and fingertip PPG data at multiple regular intervals after treadmill exercise, computes the PTT and translates it to the SBP and DBP values. [8] utilizes the ECG and PPG signals from MIMIC-II dataset, pre-processes them, extracts various physiological features, does dimensionality reduction, and implements a number of machine learning regression models to estimate SBP, MABP and DBP. [30] adopts a rather different approach for CLBP estimation whereby authors utilize data from the PPG, pressure, and height sensors (accelerometers), exploiting the fact that movement of the arm changes the natural hydrostatic pressure.
Deep learning-based methods: [31] utilizes the PPG and ABP signals corresponding to 20 subjects from the MIMIC II dataset, and implements a convolutional neural network (CNN) for feature extraction, followed by an long short-term memory (LSTM) network for CLBP prediction. [32] considers 3 subjects from MIMIC II dataset, and utilizes a CNN-LSTM model for ABP waveform synthesis. Paviglianiti et al. [33] utilize both PPG and ECG signals form the MIMIC dataset, and implement a number of DL models, i.e., ResNet, LSTM, WavNet, and more, for CLBP prediction. Brophy et al. [34] utilize PPG signals from the UCI dataset, implement a number of generative adversarial networks (GAN) models in order to realize a federated learning approach for ABP waveform synthesis. They evaluate the performance of their GAN-based federated learning model on a new unseen dataset, the University of Queensland vital signs dataset. Finally, [35] utilizes a CNN, while [36] implements a feedforward neural network for CLBP estimation.
III The UCI Dataset & Data Pre-processing
III-A The UCI public dataset on CLBP estimation
[…]
III-B Data Preprocessing
In order to prepare the data for our DL models, we implemented the following data preprocessing steps on the PPG and ABP signals. 1) We began by removing out-of-band noise from the data using a Butterworth low-pass filter with a cut-off frequency of 10 Hz. 2) We then proceeded to remove the residual artifacts manually (by discarding the anomalous samples). 3) Next, we removed the baseline from the data using the Wavelet filtering. 4) We then normalized the data using z-score normalization. 5) Further, the process of ABP waveform synthesis from PPG waveform required beat-level synchronization between the PPG and ABP signals. To this end, we utilized time shifting approach to align the systolic peaks of the PPG signal with the systolic peaks of the ABP signal. 6) We then did segmentation (with a stride of 0) to construct a number of segments from each PPG and ABP recording, in order to increase the size of the dataset. Note that each segment consists of two cardiac cycles.
IV Methodology
Keeping in mind that the problem at hand—ABP waveform synthesis from PPG waveform—is a sequence-to-sequence translation (i.e., regression) problem, we have implemented two novel deep learning methods using Tensorflow and Keras frameworks in Python: 1) a transformer model, 2) a linear/non-linear frequency-domain regression model. Once the ABP waveform is synthesized, we sample its maximum and minimum values which yield the systolic and diastolic blood pressure, respectively. Below, we discuss the two methods in detail, one by one.
IV-A Transformer model for ABP waveform synthesis
The motivation for implementing a transformer model for ABP waveform synthesis from PPG waveform comes from the following: i) both the PPG and the ABP waveforms represent the time-series data, ii) each of the two time-series consists of an ordered set of data points that capture the temporal dynamics of the heart, i.e., the heart physiology, at different time-scales and on its own, iii) there is one-to-one correspondence (relationship) between the (ordered) data points (at the same instant) in the PPG and ABP time-series, due to the fact that they both capture the same phenomenon (of heart physiology), though independently. Keeping in mind all these strong analogies with the natural language processing (NLP) problems (e.g., language translation, paraphrasing, etc.), we solve the ABP waveform synthesis problem using the transformer model—a powerful tool for solving a range of NLP problems.
Our transformer model implementation consists of the following discrete components: positional encoding, multihead self-attention, feedforward network, layer normalization, dropout, and a dense output layer. Fig. 1 shows the architecture of the transformer implemented in this work in detail.
Figure 1: Architecture of Transformer model for ABP waveform synthesis.
At this point, it is imperative to revisit the individual building blocks of the transformer model to see how each component helps solve the problem at hand. 1) First, the positional encoding (often implemented by means of sine and cosine functions) keeps the transformer informed about the order and position of data points (i.e., the notion of time) in each of the two (i.e., PPG and ABP) sequences. 2) Next, the multi-head self-attention mechanism, as the name implies, focuses on capturing relationship between every pair of data points (by means of attention scores) in both the PPG and ABP sequences, using multiple heads that attend to different parts of the PPG and ABP sequences in order to capture the short term and the long term dependencies within the PPG and ABP signals. 3) Next, the layer Normalization reduces the internal covariate shift, prevents the issue of vanishing or exploding gradients, helps prevent overfitting, all by normalizing the activations. 4) The dropout regularization layer randomly sets a fraction of the values to zero during training, which helps to prevent overfitting. 5) Element-wise addition between the original input and the output of the dropout layer is performed in the next block. This residual connection allows information to bypass the self-attention and normalization layers, helping the model retain valuable information from the original input. 6) Next, the feedforward network applies a set of linear transformations and non-linear activations to the input data, allowing the model to learn complex patterns and relationships. 7) Finally, the dense layer produces the output of the model.
IV-B Linear/non-linear frequency-domain regression model for ABP waveform synthesis
In this method, we follow a two-step approach for ABP waveform synthesis. During the training phase, the two steps are implemented as follows:
  • Frequency-domain representation of pre-processed data: Firstly, we obtain the frequency-domain representation of the pre-processed data, for both signals of interest, i.e., PPG and ABP signals. The motivation for this comes from the fact that both PPG and ABP signals are quasi-periodic due to the rhythmic activity of the heart. Specifically, we consider one PPG-ABP segment-pair corresponding to a single cardiac cycle of duration $Q^{\prime}$ samples, at a time. We then compute the discrete cosine transform (DCT) (type-II) of each of the PPG and ABP segments in the single PPG-ABP segment-pair. Further, it is well-known that the PPG and ABP signals are sparse in frequency domain (with only a few low-frequency components with significant energy). This allows us to have a compact representation of each of the two segments in the given PPG-ABP segment-pair by retaining first $Q_{X}$ ($Q_{Y}$) significant DCT coefficients of the PPG (ABP) segment only. This in turn helps in reducing the computationally complexity of our linear/non-linear (L/NL) regression model by reducing the dimension of the feature vector it takes as input. Note that we zero-pad both DCT vectors of size $Q_{X}$ and $Q_{Y}$ to resize them to $Q$ in order to do the L/NL regression in the next step (where $Q$ represents the duration of a cardiac cycle, in samples).
[…]
Next, the testing phase. We compute DCT (type-II) of a previously unseen size-$Q^{\prime}$ PPG segment, resize it to $Q_{X}$ and then $Q$. We then feed the size-$Q$ DCT vector (of PPG) to the L/NL frequency-domain regression model which outputs a size-$Q$ DCT vector (of ABP). Then, ABP waveform is synthesized by taking the inverse DCT (I-DCT) of the output of the L/NL regression model. Fig. 2 provides a pictorial overview of the proposed FD L/NL regression model.
Figure 2: Architecture of FD L/NL regression model for ABP waveform synthesis.
V Results and Discussion
Performance Metrics: We have utilized the following two metrics to evaluate the performance of our DL models for CLBP estimation: mean absolute error (MAE), and root mean squared error (RMSE), where MAE = $\frac{1}{N}\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|$, and $\text{RMSE}(y,\hat{y})=\sqrt{\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{N}}$. Here, $N$ represents the batch size, $y$ represents the reference value/ground truth, and $\hat{y}$ represents the model prediction.
V-A Performance evaluation of Transformer model
Setting of hyperparameters: We used Adam optimizer, and used a learning rate of 1e-3. Further, we set segment size to 2 seconds, i.e., 250 samples. For training and evaluation of transformer model, we utilized an 80-20 train-test split. In multi-head attention layer, we set number of heads to 8, with the size of each attention head for query and key being 64. Next, the feedforward network consisted of 64 neurons and a relu activation function. The number of transformer layers was varied between 7-15. For layer normalization, we set epsilon to 1e-6. Finally, we set the dropout ratio to 0.1.
Fig. 3 is a qualitative result that showcases two examples where the transformer model synthesizes the ABP signal from a PPG signal. We make the following observations: 1) It is clear that the morphology of the two synthetic ABP waveforms constructed by the transformer model (on the right) is quite similar to the morphology of the reference ABP waveforms (on the left). 2) We note that the synthetic ABP waveform is shifted in amplitude, compared to the reference ABP waveform. This is the root cause of the error when we measure the SBP and DBP later by sampling the maximum and minimum point of the synthetic ABP waveform.
Now, the quantitative results. The lowest MAE (for ABP waveform synthesis) achieved by the transformer model is 9.10, with 15 layers. That is, the transformer model’s performance increases with increase in number of layers, and reaches a maximum for 15 layers (as further increasing the depth does not result in significant improvements). Further, the transformer model achieved an MAE of 19 and 9, for DBP and SBP, respectively.
Figure 3: Two examples of ABP waveform synthesis by our transformer model.
V-B Performance evaluation of FD L/NL Regression model
Setting of hyperparameters: For ridge regression, we set $\alpha=0.5$. The rest of the hyperparameters remain the same.
Fig. 4 is again a qualitative result that demonstrates by means of two examples the quality of the ABP signal synthesized by the FD L/NL regression model. It is again evident that the morphology of the two synthetic ABP waveforms produced by the FD L/NL regression model (on the right) and the morphology of the reference ABP waveforms (on the left) are highly similar. Also, the synthetic ABP waveform is again shifted in amplitude, compared to the reference ABP waveform. But luckily, this time this shift in amplitude of the synthetic ABP is less pronounced. This helps us get a better MAE performance when we measure the SBP and DBP.
For our FD L/NL regression model, after converting the PPG signal to frequency-domain by means of the DCT, we have implemented ridge regression for linear regression purpose, and a feedforward neural network (FFNN) and a recurrent neural network (RNN) for non-linear regression purposes. The resulting MAE performance (for ABP waveform synthesis) is summarized in Table I. We learn from Table I that the ridge regression outperforms the FFNN and RNN by a large margin, which indicates that the relationship between the PPG signal and the ABP signal is perhaps linear.
Finally, our FD L/NL regression model (with ridge regression) achieves an MAE of 8.01 mmHg for DBP and an MAE of 11.87 mmHg for SBP.
TABLE I: Performance of different architectures for our FD L/NL regression model
FD L/NL regression model MAE
DCT + Ridge Regression 8.1
DCT + RNN 17.23
DCT + FFNN 20.24
Figure 4: Two examples of ABP waveform synthesis by our FD L/NL regression model.
V-C Performance comparison with related work
Table II provides a detailed performance comparison of our work with the selected related works by providing a quick summary of different deep learning models used for CLBP estimation, their corresponding datasets, and the performance achieved by each model. Table II efficiently highlights the variability in the performance of the related works, which is in turn due to the change of data distribution across datasets, variable sizes of the datasets, model architectures, etc. Our first approach, i.e., the transformer based model achieved an MAE of 9.10 for ABP waveform synthesis which is at par with the related works. However, the transformer model achieved a relatively large MAE of 19, for DBP estimation, which suggests that our model is well-trained for ABP synthesis and SBP estimation only. Our 2nd approach, i.e., the FD L/NL regression model achieved a MAE of 8.01 for DBP and 11.87 for SBP, which is again at par with the performance reported by the related works.
TABLE II: Performance comparison of our work with the related work: SBP MAE and DBP MAE are in mmHg. $N$ represents number of subjects. * represents RMSE in mmHg.
Work
Dataset
$N$
Algorithm
SBP
DBP
Tazarv et al. [31]
MIMIC II
20
CNN-LSTM
3.70
2.02
Paviglianiti et al. [33]
MIMIC
40
ResNet, LSTM, WavNet, ResNet+LS-TM
4.118
2.228
Treebup-achatsa-kul et al.
UCI
812
CAN
7.14*
6.08*
Slapnicar et al.
MIMIC III
510
Spectrotem-poral ResNet
9.43
6.88
Wang et al.
UCI
348
Pretrained AlexNet, Inception-V3, VGG-19
6.17
3.16
Aguirre et al.
MIMIC
1131
RNN Encoder-Decoder + attention
14.39
14.39
Our work (transformer)
UCI
209
Transformer
19.0
9.0
Our work (FD L/NL regression)
UCI
209
DCT+L/NL Regression
11.87
8.01
TABLE III: Performance comparison of our work with related work according to the BHS criterion
Work Grade Cumulative Error
$\leq 5\%$ $\leq 10\%$ $\leq 15\%$
BHS A 60% 85% 95%
B 50 75 90
C 40 65 85
N. Ibtehaz et. al [18] A 87.4 95.2 97.7%
Tazarv et. al [31] A 75 92 96%
Li Y-H et. al B 59.46 79.97 88.45%
Kauchee et. al [8] C 44.7 71.6 86.7%
Yang et. al C 42.1 94.7 100%
Our work (FD L/NL regression) B 51.69 76.79 100%
Our work (transformer) C 36.47 71.98 100%
Finally, Table III outlines the performance grades of our transformer model and FD L/NL regression model, according to the BHS grading criterion. Specifically, Table III shows the cumulative percentages (CP) of the error achieved by the two models on the UCI test data. We can see that our transformer model achieves the grade C, while our FD L/NL regression model achieves the grade B. For the sake of completeness, Table III also summarizes the BHS performance grades obtained by the selected related works, for comparison purposes. We also want to highlight that nearly all the works suffer from the issue of lack of generalization, i.e., as the number of subjects increases, the MAE of the SBP and DBP increases, while the BHS grade tends to fall. On a side note, it is also worth mentioning that both of our models models fulfill the AAMI criterion of utilizing more than 85 subjects for training purposes.
VI Conclusion
We evaluated two novel purpose-built DL models on public UCI dataset for CLBP estimation using a single-site PPG signal. The transformer model (the FD L/NL regression model) achieved an MAE of 19 and 9 (11.87 and 8.01), for DBP and SBP, respectively. Our FD L/NL regression model also fulfilled the AAMI criterion of utilizing data from more than 85 subjects, and achieved grade B by the BHS criterion.
Now, some parting thoughts. The UCI dataset consists of data from patients in intensive care units, who are often under the effects of drugs and other medical interventions. Thus, it is difficult to generalize the results of studies–that use the UCI dataset–to other populations, such as healthy adults. Thus, there is a need to evaluate our DL models on larger and diverse datasets in order to increase their generalization capability.